OpenStudy (anonymous):

please help! But also tell me how you got your answer! <3 <--- I only give metals of you follow this rule... What value of n solves the equation? 3n = 1 over 81 n=

OpenStudy (vlery):

just divide 3 into (1/81)

OpenStudy (jack1):

3n = 1 over 81 3n = 1/ 81 81 x 3n = 1 243n = 1 n = 1/243

OpenStudy (jack1):

=0.004115... ish

OpenStudy (anonymous):

 or